You randomly select two letters from eight tiles that spell G.E.O.M.E.T.R.Y.
replacing the first before you select the second,what is the probability that you will select two E's

Answer by AnlytcPhil(1806)

You randomly select two letters from eight tiles that spell G.E.O.M.E.T.R.Y.
replacing the first before you select the second,what is the probability that
you will select two E's

P(selecting E the 1st time) = 2 out of 8 = 2/8 = 1/4
P(selecting E the 2nd time) = 2 out of 8 = 2/8 = 1/4

Since we are replacing the letter chosen before choosing the second letter,
the events are independent and we can muiltiply the probabilities:

P(selecting E 1st time AND selecting E 2nd time) = 

P(selecting E 1st time) × P(selecting E 2nd time) = (1/4)×(1/4) = 1/16

Edwin
AnlytcPhil@aol.com